# The Lazy Student’s Dream: ChatGPT Passing an Engineering Course on Its Own

### v1

@@ Abstract
thanks: This work was supported by the Grainger College of Engineering at the University of Illinois Urbana-Champaign through the Grants for Advancement of Teaching in Engineering program.
@@ 1 Introduction
The rapid emergence of Large Language Models (LLMs) in educational settings (Xu et al., 2024) has sparked substantial discussion about their impact on academic integrity and learning outcomes. While numerous studies have examined LLMs’ capabilities in specific tasks (Jauhiainen and Guerra, 2024) or isolated assignments (Padiyath et al., 2024; Vadaparty et al., 2024), there remains a critical gap in our understanding of their performance across an entire technical course. This work presents a systematic investigation of LLM performance across a complete undergraduate Aerospace Control Systems course (AE 353) at the University of Illinois Urbana-Champaign (UIUC), with the goal of quantifying the LLM’s capabilities in handling diverse engineering coursework with no expert interaction and informing future course design strategies.

### v2

@@ 1 Introduction
The rapid emergence of Large Language Models (LLMs) in educational settings (Xu et al., 2024) has sparked substantial discussion about their impact on academic integrity and learning outcomes. While numerous studies have examined LLMs’ capabilities in specific tasks (Jauhiainen and Guerra, 2024) or isolated assignments (Padiyath et al., 2024), there remains a critical gap in our understanding of their performance across an entire technical course. This work presents a systematic investigation of LLM performance across a complete undergraduate Aerospace Control Systems course (AE 353) at the University of Illinois Urbana-Champaign (UIUC), with the goal of quantifying the LLM’s capabilities in handling diverse engineering coursework with no expert interaction and informing future course design strategies.
@@ 6 Conclusion
Acknowledgments
This work was supported by the Grants for Advancement of Teaching in Engineering program at the Grainger College of Engineering, University of Illinois Urbana-Champaign. We gratefully acknowledge Prof. Timothy Bretl for providing the course materials, assignments, projects, and PrairieLearn infrastructure foundational to this study. We also thank Grayson Schaer for developing the project environments and Pranay Thangeda for contributing to the PrairieLearn question bank and lecture materials. Additionally, we extend appreciation to all others whose efforts supported this research.